OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 36 inches. Find the width of the rectangle. 6 in. 11 in. 7 in.

Parth (parthkohli):

The perimeter of a rectangle is \(2 \times (length + width)\)

OpenStudy (anonymous):

ok

Parth (parthkohli):

We can take width as x, then length is 4 inches more, so: Width = x Length = x + 4 Equation we get: \( \color{Black}{\Rightarrow 2(x + x + 4) = 36 }\)

Parth (parthkohli):

Next step would be to?

OpenStudy (anonymous):

distrbute

Parth (parthkohli):

No need to distribute. Divide both sides by 2 :)

OpenStudy (anonymous):

x+x+4=18

Parth (parthkohli):

Yes, now combine like terms.

OpenStudy (anonymous):

2x+4=18

Parth (parthkohli):

Next step would be to?

OpenStudy (anonymous):

2x=14

Parth (parthkohli):

Next step :)

Parth (parthkohli):

You divide both sides by 2.

OpenStudy (anonymous):

ya was thinking was still 18

OpenStudy (anonymous):

x=7

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

thanks parth

Parth (parthkohli):

:)